A faucet can fill a sink in 5 minutes while a drain will empty the same sink in 8 minutes. If the faucet is turned on and the drain is left open, how long will it take to fill the sink?
Answer by stanbon(75887)

A faucet can fill a sink in 5 minutes while a drain will empty the same sink in 8 minutes. If the faucet is turned on and the drain is left open, how long will it take to fill the sink?
------
faucet rate = 1/5 job/min
drain rate = 1/8 job/min
together rate = 1/x job/min
-------
Equation:
faucet rate - drain rate = 1/x
1/5 - 1/8 = 1/x
---
8x - 5x = 40
3x = 40
x = 13 1/3 minutes (time to fill the sink)
===============
Cheers,
Stan H.